# Inverse Concave-Utility Reinforcement Learning is Inverse Game Theory

### v1

@@ 1 Introduction
Imagine that we are trying to infer the reward function of a human who has information-processing bounds, which is a form of bounded rationality due to cognitive constraints [26, 27, 28, 29]. This setting is often modelled as information-limited decision-making that is equivalent to the cautious / human-regularised RL formulation in table LABEL:tbl:table [30, 31, 32]. Here, $d^{\rho}$ represents a prior, low mental-effort policy for the human, and they pay a cost for deviating from it proportional to the KL divergence. The figure 1 presents an illustrative example, where a captain must navigate to a port  [^1] The shark, ship, and port icons created by Gilberto, hartadidesign, and Darwin Mulya from the Noun Project. . The captain’s true reward function is such that each time-step has a small cost, crossing into dangerous waters (represented as sharks) have a high cost, and the rewards at the ports induce a preference ordering so that the port A is preferred the most, and C the least. The environment is not stochastic, thus actions are always executed correctly. Here, the obviously optimal policy for the captain is to deterministically follow the golden path. However, being under a heavy cognitive burden, the captain knows they are prone to making errors. Thus instead, they follow a more forgiving path to port B, which leaves room for error. In similar episodes, they do make mistakes, and we end up observing a dataset of demonstrations within the green-shaded region. This is not the same as having a stochastic environment, since another captain might successfully execute the golden path in the same environment, and we already know the dynamics are deterministic. Here, even in the limit of infinite demonstrations, we cannot infer the true reward function with standard IRL. As we prove in lemma 7, the problem is more severe. We may be unable to infer any feasible state-action reward function from the optimal policy of a CURL problem by applying the standard IRL formulation.
@@ 3 Inverse Game Theory for CURL
Both inverse RL and inverse game theory share the same objective: to find a reward function that would rationalize the observed behaviour of agents. Since the lemma 4 states there cannot be a stationary reward function $R(s,a)$ in general for CURL problems, the standard formulation of IRL does not apply any more. In particular, the definition of the feasible rewards set $\mathcal{R}_{\mathbb{B}}$ is problematic for the CURL setting. It assumes that $\pi^{E}$ is optimising for the standard RL objective with stationary reward functions that only depend on $(s,a)$ and the set condition relies on the advantage function. However, as we know from the lemma 4, there are optimal CURL policies that cannot be induced by such reward functions via the linear objective. Let $\mathbb{C}=(\bar{M},\pi^{E})$ denote an I-CURL problem where $\pi^{E}$ is the optimal policy for a concave utility function $F$. The following lemma states the impossibility of applying standard IRL to $\mathbb{C}$ in general.
@@
There exists an I-CURL problem $\mathbb{C}=(\bar{M},\pi^{E})$ such that the $\mathcal{R}_{\mathbb{C}}=\emptyset,$ where $\mathcal{R}_{\mathbb{C}}$ is the feasible reward set for a standard IRL problem, as defined in definition 6.
Proof. Let $(M,F)$ be a CURL problem with expert policy $\pi^{E}\in\operatorname*{argmax}_{\pi}F(d^{\pi}).$ The lemma 4 states that there may be no reward function $R\in\mathbb{R}^{S\times A}$ such that $\pi^{E}\in\operatorname*{argmax}_{\pi}\langle d^{\pi},R\rangle.$ This is equivalent to saying there may be no $R$ such that $A^{\bar{M}\cup R}_{\pi^{E}}(s,a)\leq 0$ everywhere.

### v2

@@ 1 Introduction
Imagine that we are trying to infer the reward function of a human who has information-processing bounds, which is a form of bounded rationality due to cognitive constraints [26, 27, 28, 29]. This setting is often modelled as information-limited decision-making that is equivalent to the cautious / human-regularised RL formulation in table LABEL:tbl:table [30, 31, 32]. Here, $d^{\rho}$ represents a prior, low mental-effort policy for the human, and they pay a cost for deviating from it proportional to the KL divergence. The figure 1 presents an illustrative example, where a captain must navigate to a port  [^1] The shark, ship, and port icons created by Gilberto, hartadidesign, and Darwin Mulya from the Noun Project. . The captain’s true reward function is such that each time-step has a small cost, crossing into dangerous waters (represented as sharks) have a high cost, and the rewards at the ports induce a preference ordering so that the port A is preferred the most, and C the least. The environment is not stochastic, thus actions are always executed correctly. Here, the obviously optimal policy for the captain is to deterministically follow the golden path. However, being under a heavy cognitive burden, the captain knows they are prone to making errors. Thus instead, they follow a more forgiving path to port B, which leaves room for error. In similar episodes, they do make mistakes, and we end up observing a dataset of demonstrations within the green-shaded region. This is not the same as having a stochastic environment, since another captain might successfully execute the golden path in the same environment, and we already know the dynamics are deterministic. Here, even in the limit of infinite demonstrations, we cannot infer the true reward function with standard IRL. We prove in lemma 7 that indeed the true reward function of a CURL problem may never be inferred via standard IRL even in the case of known expert policy.
@@ 3 Inverse Game Theory for CURL
Both inverse RL and inverse game theory share the same objective: to find a reward function that would rationalize the observed behaviour of agents. Since the lemma 4 states there cannot be a stationary reward function $R(s,a)$ in general for CURL problems, the standard formulation of IRL does not apply any more. In particular, the definition of the feasible rewards set $\mathcal{R}_{\mathbb{B}}$ is problematic for the CURL setting. It assumes that $\pi^{E}$ is optimising for the standard RL objective with stationary reward functions that only depend on $(s,a)$ and the set condition relies on the advantage function. However, as we know from the lemma 4, there are optimal CURL policies that cannot be induced by such reward functions via the linear objective. Let $\mathbb{C}=(\bar{M},\pi^{E})$ denote an I-CURL problem where $\pi^{E}$ is the optimal policy for a concave utility function $F$. The following lemma states that even in the exact case with true knowledge of $\pi^{E}$, the true reward function may be impossible to infer with standard IRL.
@@
There exists an MDP $M$ with concave utility $F$ and reward function $r^{*}$ such that, given its I-CURL problem $\mathbb{C}=(\bar{M},\pi^{E})$, $r^{*}\notin\mathcal{R}_{\mathbb{C}}$ where $\mathcal{R}_{\mathbb{C}}$ is the feasible reward set for a standard IRL problem, as defined in definition 6.
Proof. Consider the gridworld shown in 1 with rewards $r^{*}(\textit{port A})=10,r^{*}(\textit{port B})=5$, $r^{*}(\textit{port C})=1$, and $r^{*}(\textit{shark})=-100.$ The reward is zero elsewhere. As described, the observed $\pi^{E}$ is the green-shaded region. The uniquely optimal policy under the standard RL objective is the deterministic golden path shown on the figure. Therefore, the advantage function of any other policy will not be $\leq 0$ everywhere. Thus, $r^{*}$ will not be in $\mathcal{R}_{\mathbb{C}}.$